# Hydrostatic and explosive -element chemical abundances of Milky Way globular clusters, halo substructures, and satellite galaxies

### v1

@@ Abstract
Galaxy: abundances ā Galaxy: Stellar populations ā Galaxy: evolution1 Introduction
@@ 2 Data
The ids for the stars belonging to Galactic globular clusters (GCs) come from the GC value-added catalogue in APOGEE DR17 (Schiavon etĀ al., 2024), and include the following GCs: Djorgā2, FSRā1758, HPā1, Lillerā1, NGCā0104, NGCā0288, NGCā0362, NGCā1851, NGCā1904, NGCā2298, NGCā2808, NGCā3201, NGCā4147, NGCā4590, NGCā5024, NGCā5053, NGCā5139, NGCā5272, NGCā5466, NGCā5634, NGCā5904, NGCā6093, NGCā6121, NGCā6144, NGCā6171, NGCā6205, NGCā6218, NGCā6229, NGCā6254, NGCā6273, NGCā6293, NGCā6304, NGCā6316, NGCā6341, NGCā6380, NGCā6388, NGCā6397, NGCā6401, NGCā6441, NGCā6517, NGCā6522, NGCā6528, NGCā6539, NGCā6540, NGCā6544, NGCā6553, NGCā6558, NGCā6569, NGCā6642, NGCā6656, NGCā6715, NGCā6717, NGCā6723, NGCā6752, NGCā6760, NGCā6809, NGCā6838, NGCā7078, NGCā7089, Palā1, Palā5, Pal6, Palā10, Rupā106, Terzanā2, Terzanā4, Terzanā5, Terzanā9, Terzanā10, Terzanā12, Tonā2, and UKSā1.
@@
In addition, we also impose the following selection criteria to ensure reliable element abundance ratio measurements for red giant stars: SNR ; ; K.
@@ 2.1 Accounting for light-element variations in globular clusters
GCs are known to host multiple-populations (MPs; Bastian & Lardo, 2018). The second generation/population type (SP) are known to host light-element variations. Typically, SP stars are enriched in certain light elements whilst simultaneously being depleted in others, and this chemical profile isnāt manifested in the first population (FP) type. For example, SP stars typically show enhanced levels of Al and depleted in Mg; the same effect happens for an enhancement in either N or Na and depletion in C and O, respectively (e.g., Schiavon etĀ al., 2017; MĆ©szĆ”ros etĀ al., 2020). Thus, in order to mitigate any possible bias in our -element analysis, we must account for these chemical abundance anomalies arising from SP stars. To do so, we employ a -means (unsupervised) clustering algorithm, in order to determine the FP GC stars that we will use in our analysis. We run this clustering procedure using the kmeans package in scikit-learn (Pedregosa etĀ al., 2011), and employing the following set of abundances for all GC stars: [Mg/Fe], [Al/Fe], [C/Fe], [N/Fe]. When running the code, we choose the following hyperparameters: n-clusters; random-state; n-initauto. We find that using this set-up, we are able to successfully split FP from SP GC stars (see FigĀ 1). We also tested using a support vector machine to assess if we could obtain a better classification, but found that the -means clustering provided better overall classification performance.
@@ 3 Hydrostatic and explosive -elements and the āhexā ratio
Overall, we find that GCs and halo substructures occupy a similar locus in all of these diagrams, which sits above the dwarf satellite galaxy sample. In [/Fe] and [/Fe], the difference between GCs and halo substructures with respect to dwarf galaxies is of dex; conversely, the difference in the hex ratio between these populations is smaller, on the order of dex. Interestingly, we find that the GCs and halo populations show a decreasing trend in [/Fe] and the hex ratio with increasing [Fe/H], but approximately a flat profile in [/Fe]. This is also the case for dwarf galaxies, albeit there being an offset with the GCs/halo substructures.
@@ 3.1 Hex ratio across age and mass ā globular clusters
FigĀ 3 shows the median hex ratio now as a function of age for the Galactic GC sample. The ages are taken from the compilation of Forbes & Bridges (2010). Our results show that there appears to be a weak decreasing trend of hex with decreasing age. However, there is a significant scatter around this relation.
@@ 3.2 Hex ratio of accreted vs in situ globular clusters
FigĀ 4 shows again the median hex ratio as a function of median [Fe/H] for Galactic GCs, but now each GC is categorised into an accreted or in situ origin based on the associations from Massari etĀ al. (2019). We find that there is no dependence on origin between the different GC subgroups. All GCs follow the same decreasing trend in hex with increasing [Fe/H]. There is one clear outlier however, NGCā7078, which is a [Fe/H]-poor ([Fe/H] ) disc GC that presents a low hex value for its metallicity (hex ).
@@ 4 Summary conclusions
Interestingly, we also find that there are some outliers in each sample studied. For example, the Sagittarius dSph (Sgr) and its core (M54) have a hex ratio that more closely follows the satellite galaxy trend (see also McWilliam etĀ al., 2013; Carlin etĀ al., 2018). This is perhaps not that surprising, given that Sgr is a satellite galaxy that has recently been accreted by the Milky Way. In contrast, the Sextans dwarf galaxy has a hex ratio that places it in the middle of the GC/halo substructure distribution in the [Fe/H]-poor end (see FigĀ 5). This result implies that Sextans likely formed from a more top-heavy IMF when compared to its satellite galaxy counterparts with approximately the same [Fe/H]. We also find that several halo substructures (Pontus, Sequoia, and Wukong) have an almost identical hex ratio (and -Fe abundance) to GES, in concordance with the hypothesis from Horta etĀ al. (2023) that these are likely different parts of the GES debris. The same insight applies to Aleph and Nyx, which show -abundances and a hex ratio that is very similar to the high- disc. In addition, there is one dwarf galaxy (Bootes I) that has a hex ratio of at a low [Fe/H] . This implies that this dwarf galaxy could have formed from a significantly more top-light IMF, although see Filion etĀ al. (2022). There is one bulge GC (Lillerā1) that also has a low hex ratio of āalbeit at a much higher [Fe/H] ā (see also Liptrott et al, in prep), and another halo GC (NGCā7078) that shows a very low hex ratio for its low metallicity ([Fe/H] )222NGCā7078 shows a low hex ratio because of its elevated explosive -Fe ratio at its given [Fe/H]..
@@ Acknowledgements
DH would like to thank Jennifer Mead and Andrew Mason for help using the bitmasks in APOGEE to identify dwarf galaxy star candidates. DH would also like to thank Sue, Alex, and Debra for their support. This work was supported by the UKRI Science and Technology Facilities Council under project 101148371 as a Marie Curie Research Fellowship.

### v3

@@ Abstract
Galaxy: abundances ā globular clusters: general ā Galaxy: evolution1 Introduction
@@ 2 Data
The ids for the stars belonging to the Galactic globular clusters (GCs) used in this work come from the GC value-added catalogue in APOGEE DR17 (Schiavon etĀ al., 2024), and include the following GCs: Djorgā2, FSRā1758, HPā1, NGCā0104, NGCā0288, NGCā0362, NGCā1851, NGCā1904, NGCā2298, NGCā2808, NGCā3201, NGCā4147, NGCā4590, NGCā5024, NGCā5053, NGCā5139, NGCā5272, NGCā5466, NGCā5904, NGCā6121, NGCā6171, NGCā6205, NGCā6218, NGCā6229, NGCā6254, NGCā6273, NGCā6293, NGCā6304, NGCā6316, NGCā6341, NGCā6380, NGCā6388, NGCā6397, NGCā6441, NGCā6522, NGCā6540, NGCā6544, NGCā6553, NGCā6558, NGCā6569, NGCā6642, NGCā6656, NGCā6715, NGCā6717, NGCā6723, NGCā6752, NGCā6760, NGCā6809, NGCā6838, NGCā7078, NGCā7089, Palomarā5, Palomarā6, Terzanā2, Terzanā4, Terzanā9. We follow the recommendation of Schiavon etĀ al. (2024) and select likely GC star candidate members by enforcing dPOS (stars withing one Jacobi radii), dRV (stars within from the GCās line-of-sight velocity distribution), and dPM (stars within from the GCās proper motion distribution).
@@
In addition, we also impose the following selection criteria to all GCs, halo substructures, and satellite galaxies to ensure reliable element abundance ratio measurements for red giant stars: SNR ; ; K.
@@ 2.1 Accounting for light-element variations in globular clusters
GCs are known to host multiple-populations (MPs; Bastian & Lardo, 2018). The second generation/population type (SP) are known to host light-element variations. Typically, SP stars are enriched in certain light elements whilst simultaneously being depleted in others, and this chemical profile isnāt manifested in the first population (FP) type. For example, SP stars typically show enhanced levels of Al and depleted in Mg; the same effect happens for an enhancement in either N or Na and depletion in C and O, respectively (e.g., Schiavon etĀ al., 2017; MĆ©szĆ”ros etĀ al., 2020). Thus, in order to mitigate any possible bias in our -element analysis, we must account for these chemical abundance anomalies arising from SP stars. To do so, we employ a -means (unsupervised) clustering algorithm, in order to determine the FP GC stars that we will use in our analysis. We run this clustering procedure using the kmeans package in scikit-learn (Pedregosa etĀ al., 2011), and employing the following set of abundances for all stars belonging to each GC: [Mg/Fe], [Al/Fe], [C/Fe], [N/Fe]. When running the code, we choose the following hyperparameters: n-clusters; random-state; n-initauto. We find that using this set-up, we are able to successfully split FP from SP GC stars (see FigĀ 1). We also tested using a support vector machine to assess if we could obtain a better classification, but found that the -means clustering provided similar classification performance.
@@ 3 Hydrostatic and explosive -elements and the āhexā ratio
Overall, we find that GCs and halo substructures occupy a similar locus in all of these diagrams, which sits above the dwarf satellite galaxy sample. In [/Fe] and [/Fe], the difference between GCs and halo substructures with respect to dwarf galaxies is of dex; conversely, the difference in the hex ratio between these populations is smaller, on the order of dex. Furthermore, we find that the GCs and halo populations show a decreasing trend in [/Fe] and the hex ratio with increasing [Fe/H], but approximately a flat profile in [/Fe]. This is also the case for dwarf galaxies, albeit there being an offset with the GCs/halo substructures.
@@ 3.1 Hex ratio across age and mass ā globular clusters
FigĀ 3 shows the median hex ratio now as a function of age for the Galactic GC sample. The ages are taken from the compilation of Forbes & Bridges (2010). Our results show that there appears to be a weak decreasing trend of hex with decreasing age222We calculate a slope of dex with increasing age and an intercept of dex.. However, there is a significant scatter around this relation.
@@ 3.2 Hex ratio of accreted vs in situ globular clusters
FigĀ 4 shows again the median hex ratio as a function of median [Fe/H] for Galactic GCs, but now each GC is categorised into an accreted (triangles) or in situ (circles) origin based on the associations from Massari etĀ al. (2019). We find that there is no dependence on origin between the different GC subgroups. All GCs follow the same decreasing trend in hex with increasing [Fe/H]. There are two clear outliers however, NGCā7078 which is a [Fe/H]-poor ([Fe/H] ) disc GC, and NGCā6341 ([Fe/H] ) that is a GSE GC. Both of these GCs present a low hex value for their metallicity (hex ).
@@ 4 Summary conclusions
Extra-Galactic studies modelling the luminosity of galaxies ranging in mass and environment have found that those with higher star formation rates tend to have formed from a more top-heavy IMF (Hoversten & Glazebrook, 2008; Meurer etĀ al., 2009; Lee etĀ al., 2009; Gunawardhana etĀ al., 2011; Cappellari etĀ al., 2012; Weidner etĀ al., 2013; Li etĀ al., 2017; De Masi etĀ al., 2018); follow up studies also suggest that the IMF possibly varies with metallicity or time (Li etĀ al., 2023; MartĆn-Navarro etĀ al., 2024; Bate, 2025). Assuming that the cause for the variation in the hex ratio at constant metallicity between the GC/halo substructures and the living satellites is due to variations in their star formation histories, and thus the high-mass end of the IMF, our results support these extra-Galactic studies.
@@
Interestingly, we also find that there are some outliers in each sample studied. For example, the Sagittarius dSph (Sgr) and its core (M54) have a hex ratio that more closely follows the satellite galaxy trend (see also McWilliam etĀ al., 2013; Carlin etĀ al., 2018). This is perhaps not that surprising, given that Sgr is a satellite galaxy that has recently been accreted by the Milky Way. In contrast, the Sextans dwarf galaxy has a hex ratio that places it in the middle of the GC/halo substructure distribution in the [Fe/H]-poor end (see FigĀ 5). This result implies that Sextans likely formed from a more top-heavy IMF when compared to its satellite galaxy counterparts with approximately the same [Fe/H]. We also find that several halo substructures (Pontus, Sequoia, and Wukong) have an almost identical hex ratio (and -Fe abundance) to GES, in concordance with the hypothesis from Horta etĀ al. (2023) that these are likely different parts of the GES debris. The same insight applies to Aleph and Nyx, which show -abundances and a hex ratio that is very similar to the high- disc. In addition, there is one dwarf galaxy (Bootes I) that has a hex ratio of at a low [Fe/H] . This implies that this dwarf galaxy could have formed from a significantly more top-light IMF, although see Filion etĀ al. (2022). There are also two GCs (NGCā6341 and NGCā7078) that show a lower hex ratio () for their low metallicity ([Fe/H] ).
@@ Acknowledgements
The authors thank the anonymous referee for their helpful report. DH would like to thank Jennifer Mead and Andrew Mason for help using the bitmasks in APOGEE to identify dwarf galaxy star candidates, and Andrea Kunder for helpful discussions. DH would also like to thank Sue, Alex, and Debra for their support. This work was supported by the UKRI Science and Technology Facilities Council under project 101148371 as a Marie Curie Research Fellowship.